40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 5 01:58:54 2014, ericq, Update, LSC, 3F RFPD RF spectra Nov52014_3fPD_DRMIspectra.zipREFL165_DRMIspectrum.pngREFL33_DRMIspectrum.png
    Reply  Wed Nov 5 11:09:44 2014, Koji, Update, LSC, 3F RFPD RF spectra 
       Reply  Wed Nov 5 22:25:42 2014, ericq, Update, LSC, 3F RFPD RF spectra REFL165_demod_DRMIspectrum.png
          Reply  Thu Nov 6 01:58:55 2014, Koji, Update, LSC, 3F RFPD RF spectra 
             Reply  Thu Nov 6 11:49:58 2014, ericq, Update, LSC, 3F RFPD RF spectra 
                Reply  Thu Nov 6 14:41:49 2014, Koij, Update, LSC, 3F RFPD RF spectra 
                   Reply  Fri Nov 7 02:21:12 2014, ericq, Update, LSC, 3F RFPD RF spectra 
                      Reply  Fri Nov 7 14:41:18 2014, ericq, Update, LSC, 3F RFPD RF spectra 165_ampSaturation.png165_HPatPD.png
                         Reply  Fri Nov 7 16:15:53 2014, Jenne, Update, LSC, 3F RFPD RF spectra 
                            Reply  Sat Nov 8 11:35:05 2014, rana, Update, LSC, 3F RFPD RF spectra 
                            Reply  Mon Nov 10 18:11:57 2014, ericq, Update, LSC, 3F RFPD RF spectra 165signals.pdf
Message ID: 10683     Entry time: Fri Nov 7 02:21:12 2014     In reply to: 10682     Reply to this: 10685
Author: ericq 
Type: Update 
Category: LSC 
Subject: 3F RFPD RF spectra 

 

After some enlightening conversation with Koji, we figured that the RF amplifier in the REFL165 chain is probably being saturated (the amp's 1dB compression is at +3dBm, has 23dB gain, and there are multiple lines above -20dBm coming out of the PD). I took a few more spectrum measurements to quantify the consequences, as well as a test with the highpass connected directly to the PD output, that should reduce the power into the amplifier. However, I am leaving everything hooked back up in its original state (and have removed all couplers and analyzers...)

I also took some DRMI sensing measurements. In the simple Michelson configuration, I took TFs of each ITMs motion to AS55Q to make sure the drives were well balanced. They were. Then, in the DRMI, I took swept sine TFs of PRCL, SRCL and differential ITM MICH motion to the Is and Qs of AS55 and all of the REFLs. I constrained the sweeps to 300Hz->2kHz; the loops have some amount of coupling so I wanted to stay out of their bandwidth. I also took a TF of the pure BS motion and BS-PRM MICH to the PDs. From these and future measurements, I hope to pursue better estimates of the sensing matrix elements of the DRMI DoFs, and perhaps the coefficients for compensating both SRCL and PRCL out of BS motion. 

I'm leaving analysis and interpretation for the daytime, and handing the IFO back to Diego...

ELOG V3.1.3-